In a camp, there was enough food to feed 2000 people on equal rations for 90 days. After 20 days, 500 more people joined the camp. Calculate the number of days that the remaining food would be used to feed the people.
Answer by ankor@dixie-net.com(22740)

In a camp, there was enough food to feed 2000 people on equal rations for 90 days.
After 20 days, 500 more people joined the camp.
Calculate the number of days that the remaining food would be used to feed the people.
:
Find how many meal-man days available
2000 * 90 = 180000
:
let d = no. of remaining days available when 500 men added
:
20(2000) + 2500d = 180000
40000 + 2500d = 180000
simplify, divide by 100
400 + 25d = 1800
25d = 1800 - 400
25d = 1400
d = 1400/25
d = 56 days of food after the increase to 2500 men